OpenStudy (anonymous):

What is the solution to the equation 21^(x + 3) = 14? Algebra 2.

OpenStudy (anonymous):

@Haseeb96 @Here_to_Help15

OpenStudy (freckles):

you can take ln( ) of both sides

OpenStudy (anonymous):

What do I do with (x+3)?

OpenStudy (freckles):

you can use power rule after taking ln( ) of both sides

OpenStudy (freckles):

\[\ln(21^{x+3})=\ln(14) \\ (x+3)\ln(21)=\ln(14)\] can you finish solving for x?

OpenStudy (anonymous):

\[x=\ln(14)/\ln(21)-3?\]